OpenStudy (superhelp101):

(x+1)/(3) I need the excluded values, set notion D and R, Interval notation D and R !!!!!!!!!!!!!!! plzzzz

OpenStudy (superhelp101):

@satellite73 any help

OpenStudy (superhelp101):

@aaronq @april115 @genius12

OpenStudy (superhelp101):

@thomaster ii need your helpppp

OpenStudy (superhelp101):

@MathLegend @mary.rojas @masri101 I need your help!!!!!!

OpenStudy (superhelp101):

plzzzzz

OpenStudy (superhelp101):

@sourwing i need your help!!!!!!!!

OpenStudy (superhelp101):

@sourwing plzzz

OpenStudy (superhelp101):

:)

OpenStudy (superhelp101):

@shamil98 @SeaTurtle113 @Bunnijjane @jim_thompson5910 @Nivi @NoelGreco

OpenStudy (superhelp101):

any help guys!! :( @brendalopez @Taylor<3sRin @justnick09 @Kilorion

OpenStudy (anonymous):

what math is this?

OpenStudy (superhelp101):

algebra 2

OpenStudy (justnick09):

3x+1

OpenStudy (justnick09):

thats all i got haha

OpenStudy (superhelp101):

lol thanks

OpenStudy (anonymous):

\[\frac{x+1}{3}\neq 3x+1\]!!

OpenStudy (anonymous):

if it is \(\frac{x+1}{3}\) there are no excluded values, since the denominator is just a number \[\frac{x+1}{3}=\frac{1}{3}x+\frac{1}{3}\]a line

OpenStudy (superhelp101):

okay so no excluded values. What about the set and interval notation? @satellite73

OpenStudy (superhelp101):

@satellite73

OpenStudy (superhelp101):

@satellite73

OpenStudy (superhelp101):

are you there @satellite73